OpenStudy (anonymous):

HELP ME PLEASE!!!!!!!!!!!!!!!!!!!!!! find the domain of the function f(x)=square root of quantity x plus eight divided by quantity x plus four times quantity x minus six Would to answer be.. x is greater than or equal to -8, x is not equal to -4, xx is not equal to 6?

OpenStudy (anonymous):

man this is hard to read like in words

OpenStudy (jackthegreatest):

do u have a description?

OpenStudy (anonymous):

\[f(x)=\frac{\sqrt{x+8}}{(x+4)(x-6)}\]

OpenStudy (jackthegreatest):

im not so sure of wat it looks like

OpenStudy (jackthegreatest):

but i think @findingemo is right

OpenStudy (anonymous):

I'm not sure either, it was given to me written out but I would assume that looks correct

OpenStudy (jackthegreatest):

oh k

OpenStudy (anonymous):

so look, can't take the square root of a negative right? so for sure \(x+8\geq 0\)

OpenStudy (anonymous):

that means \(x\geq -8\) also the denominator can't be zero either because that would make it not a number